a man of mass50 kg is standing on a 100kg plank kept on a fictionless horizontal floor. initially both are at rest if man starst walking on the plank with speed 6m/s towards right relative to the plank then amount of muscle energy spent by the man is!!????

Grade:

2 Answers

Chetan Mandayam Nayakar
13 years ago

from principle of conservation of momentum, man and plank move in opposite directions w.r.t. floor with speeds v and u respectively. conservation of momentum implies 50v=100u,v=2u, v+u=6, therefore u=2m/s and v= 4m/s,

Energy= (50/2)42 +(100/2)22 =600 Joule

'Win exciting gifts by answering the questions on Discussion Forum.

Gaurav
9 years ago
Hello Student
Number of Photons emitted , n = 2.5 x 1015
Duration of Source , T = 2 X 10-9 s
Energy of Source . E = nhv = nh/T = 8.3 X 10-10 J